the measurements of two complementary angles are in the ratio 2:13.Find the measure of the smaller angle.
Answer by stanbon(75887)

the measurements of two complementary angles are in the ratio 2:13.Find the measure of the smaller angle.
----------------------
Let the 2 angles be "x" and "90-x":
----------
x/(90-x) = 2/13
----
13x = 2(90-x)
---
13x = 180-2x
15x = 180
x = 12 degrees (The smaller angle)
=====================================
Cheers,
stan H.
================